The ten's digit of a two digit number is 2 less than the unit's digit. The value of the number is 7 times the unit's digit. Find the digits and the value of the number.
Remember the value of a two digit number is 10 times the ten digit (t), plus the digit (u).
I tried:
let u = unit digit
u-2 = tens digit (t)
value = 10t + u = 7(u + u-2)
10(u+2)+u =7u+7u-14
10u+20+u = 14u -14
11u-14u = -14-20
-3u = 6
u = -2
incorrect

Answer by Paul(988)

10-digit = x-2
Unit digit = x
EQUATION:
10(x-2)+x=7(x)
10x-20+x=7x
-4x=-20
x=5
5-2=3
Hence, the 10-digit is 3 and the unit digit is 5 the number is 35.
Paul.